Do anybody know the answer and solution I have tried several times to get the answer.



Find all solutions to the equation.

10x2 = 640

only x = 8
x = 8 or x = 64
x = 0 or x = 64
x = 8 or x = −8
only x = 64

Answer by checkley77(12844)


10x^2 = 640
x^2=640/10
x=64x=sqrt64
x=+/-8 ans.
PROOF:
10(8^2)=640
10*64=640
640=640